OpenStudy (anonymous):

will medal anyone who can help, & fan! Please help, ALGEBRA 2.

OpenStudy (anonymous):

\[5\sqrt{32}-7\sqrt{8}\] simplify

OpenStudy (anonymous):

4sqrt2

OpenStudy (anonymous):

@everlearner can you help me with one more?

OpenStudy (anonymous):

yeah bring it on

OpenStudy (anonymous):

u have got a question ??

OpenStudy (anonymous):

@everlearner you're awesome, okay. simplify this equation: \[2\sqrt{15}xy ^{3}\times3\sqrt{30}x ^{3}y ^{2}\]

OpenStudy (anonymous):

there ya go. thanks so much. @everlearner

OpenStudy (anonymous):

90sqrt2*x^4y^5

OpenStudy (anonymous):

teach me to write the expression u write .. i dont know how to :

OpenStudy (anonymous):

when you go to write a response, it has an equation button on the bottom left corner, hit it, & it will give you the space to write one. @everlearner

OpenStudy (anonymous):

@everlearner if you know the answer to this one, can explain how you got it?\[(-64)^{1/3}\] just simplify